OpenStudy (anonymous):

2x-y=5 9x-4y=25 Use the method of elimination to solve the system of linear equations

OpenStudy (unklerhaukus):

eliminate a variable by solving the first equation for y

OpenStudy (unklerhaukus):

y= ...

OpenStudy (unklerhaukus):

can you do that ? @Avon

OpenStudy (anonymous):

nope help me out @UnkleRhaukus

OpenStudy (anonymous):

is it y=30

OpenStudy (unklerhaukus):

2x-y=5 2x-5=y y=2x-5

OpenStudy (unklerhaukus):

now substitute this into the second equation 9x-4y=25 9x-4(2x-5)=25 solve this for x

OpenStudy (anonymous):

18x-8-20=25

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

I can help you @Avon

OpenStudy (anonymous):

Okay, so we want to use elimination right?

OpenStudy (anonymous):

so take the first equation and multiply it by -4 because we want to eliminate the y terms. You get: -8x+4y= -20 9x-4y=25

OpenStudy (anonymous):

okay. so add these together. 1x=-5 x=5

OpenStudy (anonymous):

Now subsitute x into the second equation and solve for y. 9(5)-4y=25 45-4y=25 -4y=-20 y=5

OpenStudy (anonymous):

Now we can check.. 2(5)-5=5 5=5 This one checks. 9(5)-4(5)=25 25=25 This one checks. (5,5) is your answer.

OpenStudy (anonymous):

Hope that helps! @Avon

OpenStudy (anonymous):

thanks @Atkinsoha

OpenStudy (anonymous):

No problem! UnkleRhaukus was doing substitution. It will be the same answer, but this problem specifically said elimination.

OpenStudy (anonymous):

U MADE IT EASY FOR ME @Atkinsoha

OpenStudy (anonymous):

No problem! Glad I could help! :) Mention me in your problem if you need any more help :)

OpenStudy (unklerhaukus):

i got the words mixed up

OpenStudy (anonymous):

It's alright @UnkleRhaukus It would have been the same answers!